1. Turned off high voltage power supplies for PZT1/2 (input PZTs) and OMC stage 1/2. They live in 1Y3 rack and AUX_OMC_NORTH rack.
2. Restored all IFO optics alignment to the position where I aligned this afternoon (for SRM, I didn't aligned it; it restored at the saved value on May 26).
3. Centered all the oplevs. They can be used for a reference for alignment change before and after the vent.
I will leave PSL mechanical shutter and green shutters closed just in case.
Some MEDM screenshots below.